OpenStudy (anonymous):

Evaluate Integral: 2sin(2x) from 0, pi/2

OpenStudy (unklerhaukus):

\[\int\limits_0^{\pi/2}2\sin(2x)\text d x\] \[=2\times\{-\frac{\cos(2x)}{2}|_0^{\pi/2}\}\]

OpenStudy (anonymous):

Did u use half or double angle formula??

OpenStudy (anonymous):

Because sin(2x)= 2sin(x)cos(x) Although, i know that antiderivative of sin is -cos.

OpenStudy (unklerhaukus):

um you might use the double angular formula if you were asked to integrate 2sin(x)cos(x) but in this question it would make the formula more complicated rater than more simple

OpenStudy (unklerhaukus):

i could have been more explicit with my integration \[2\int\sin(2x)\text dx\] let 2x=u x=u/2 dx=du/2 \[2\int\sin(u)\frac{\text d u}{2}\]

OpenStudy (anonymous):

U substitution will be the death of me, i swear. Thank you!! I get it now

OpenStudy (unklerhaukus):

i really struggled with t when i was first learning it by it is clear to me now, do you have any more question./s

OpenStudy (anonymous):

I have one on arc lengths = )

OpenStudy (unklerhaukus):

agh not my specialty

OpenStudy (anonymous):

None of this stuff is my specialty,lol. Thanks alot!!!

OpenStudy (anonymous):

Oh, considering how Calc 2 has many formulas isthere any special way u have memorized them??

OpenStudy (unklerhaukus):

well i dont remember many of trig formulas but i can usually look them up or work them out if i have to